« Project Status: Tuesday 1 January (20:30 UTC) | Main | Project Status: 3 January 2008 2030UTC »

Project Status: 2 January 2008 2030UTC

IOP #4: Big event still looks likely for Thursday to Sunday for the HMT area.

Weather Status: HMT and NWS forecasters indicate a high probability of a big event with precipitation amounts of 5-9 inches over the next 5 days. Precursors show that moisture and moisture flux are anomalously high. The event will unfold in 3 stages: Stage 1 begins Thursday evening (4 Jan 03UTC) with precipitation onset. HYDROX radar site will be right at melting level at initial stages. Stage 2 brings in a more abundant moisture surge and strong winds that will bring rain to the radar site as warm advection elevates the freezing level until Friday evening (5 Jan 03UTC). At that time a cold frontal passage occurs changing precip over to snow, likely to be very heavy. Precipitation lightens up Saturday but no clean break is evident. Stage 3 begins Sunday morning with another round of heavy snow likely.

Summary: Event due to begin near 4 Jan 03UTC, ends Sunday evening (~ 7 Jan 06UTC)

Observation Status: One radar crew member is on site doing some adjustments today and will be joined by second crew member tonight. Both will deploy to radar Thursday to prepare for operations. Today we saw little reason to modify the plannedĀ 4 Jan 00UTC start. Sondes at SLH, OAK and RNO were planned to commence at 3 Jan 20UTC, and again little was seen today to modify this start time. Ending times for both look to be Sunday evening, but will try to get more precise in coordination calls this week. Chico profiler is still down, but an expert is enroute to troubleshoot the problem.

Decision: Sondes commence 3 Jan 20UTC (12L) 4 hour intervals; Radar goes operational 4 Jan 00UTC (16L)....Radar crew should set up for a prolonged stay (to Monday) as snow will be heavy and closure of I-80 likely. Also winds will be very strong at the site.

Product Status: High resolution models running; being ingested by ESRL ALPS workstation; not making it to web server. Crew is troubleshooting this data transfer.

McGinley, GSD

TrackBack

TrackBack URL for this entry:
http://www.etl.noaa.gov/cgi-bin/mt/mt-tb.cgi/537